OpenStudy (anonymous):

im new can someone please help in this question please.Molly is having her birthday party at Pizza Party Palace. The costs involved are listed below. Pizza Party Palace requires a 15% deposit prior to the date of the event. If the deposit has been paid, how much is still owed for Molly's party? Party Prices • Food & Drink $300 • Game Tokens $125 • Other $85

OpenStudy (elonasushchik):

Add all of them together, multiple by .15. You get 72, subtract 70 to from what you got when you added all of them together which is $408

OpenStudy (anonymous):

i forgot to put the answer here they are . $76.50 $357 $153 $433.50

OpenStudy (anonymous):

So 300+125+85=510 510*.15=76.5 510-76.5=433.5

OpenStudy (anonymous):

i need more help legendarysadist

OpenStudy (anonymous):

thanks guys how do i give people medals

OpenStudy (anonymous):

I believe it is the "Best response' option that should be located next to the medals.

OpenStudy (anonymous):

thanks can you just answer this question for me please. . Nathan took his car in for service and repairs. He had a coupon for 10% off. Original Prices • Labor $212.40 • Parts $390.60 • Other $75.18 Which amount is closest to Nathan's total costs after the 10% discount and including 7% sales tax? [Assume tax is applied after the discount is applied.] $610 $563 $568 $653

OpenStudy (anonymous):

legendary sadist you didn't help me with question i was going to ask can you?

OpenStudy (anonymous):

212.4+390.6+75.18=678.18 678.18*.1=67.818 678.18-67.818=610.362 610.362*.07=42.73 610.362+42.73=653 Answer is 653